OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = -1/5 and point (5, -2). A. y + 2 = -(1/5)(x - 5) B. y + 5 = -(1/5)(x - 2) C. y – 5 = (1/5)(x + 2) D. y – 5 = (1/5)(x - 2)

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) slope(m) = -1/5 (5,-2) x1 = 5 and y1 = -2 now sub and tell me what you get

OpenStudy (texaschic101):

do not solve for y, just sub in your stuff

OpenStudy (anonymous):

mind explaining to me how to do this step by step? haven't done this type of math in 3 years

OpenStudy (texaschic101):

sure... y - y1 = m(x - x1) subbing in our info gives us : y - (-2) = -1/5(x - 5) -- but look at your left side...change it because of the signs y + 2 = -1/5(x - 5) does this make sense to you ?

OpenStudy (texaschic101):

do you have any questions ?